OpenStudy (anonymous):

Determine the number and type of complex solutions and possible real solutions for each of the following problems. 1) 2x2 + 5x + 3 = 0 2) 4x3 – 12x + 9 = 0 3) 2x4 + x2 – x + 6 = 0

OpenStudy (jack1):

2x^2 + 5x + 3 = 0 (x+1) (2 x+3) = 0 x = -1 or x = -3/2

OpenStudy (jack1):

do you understand why?

OpenStudy (anonymous):

I'm still confused on the topic....

OpenStudy (jack1):

ok, this is what's called a quadratic equation 2x^2 + 5x + 3 = 0 there are generally 2 ways to solve for the easy ones: factorization, so breaking it up into (this) times (that) = 0 in this case, it can be broken down to (x+1) (2x+3) = 0 its the reverse of the FOIL technique so now u have (this) times (that) = 0 (x+1)times (2x+3) = 0 so to get 0, either (x+1) = 0 or (2x+3) = 0

OpenStudy (jack1):

does that make sense so far?

OpenStudy (jack1):

so does the above make sense @menosleep ?